# References.bib

### v7

@@ 6.6 $\prod$-product definition step 3: $\theta(n)$
Table Tbl. 6.6.2: Index terms of the products in Tbl. 6.5.1, highlighting signs and exponents
It is evident that the sign of $k$, as well as the exponent, is represented by the same function. That function has the output pattern $[-1,1,1]$ and a F.F. that matches this requirement is $\theta(n)$:
@@
Table Tbl. 6.6.3: Index terms of the products in Tbl. 6.5.1, unified
@@
Table Tbl. 6.6.4: Roman factorial as a rising or falling product (generalized)
This result concludes Section 6, as well as the first part of this study.
@@
These alternative $\prod$-products would arise if we chose a different approach in Subsections 6.2 and 6.3, when we chose to multiply the existing products by $n$ instead of altering their upper limits.
@@ 7.1 Summary
  • Section 2: We analyzed the factorial and presented the Roman factorial.
  • Section 3: We introduced a set of 5 foundational functions (F.F.).
  • Section 4: We employed the F.F. shown in the previous section to rewrite the Roman factorial definition concisely, in its original formulation as well as its recursive form.
  • Section 5: We introduced another set of 5 F.F.
  • Section 6: We found non-recursive $\prod$-product definitions to describe the values outputted by the Roman factorial. The expressions are either a rising or a falling product, split into two cases each (positive and negative integers). Additionally, we used all F.F. to unify the cases of these definitions, achieving universality.
In Sections 3 and 5 we defined the following foundational functions:
$\delta(n)$ $=\lfloor n\rfloor+0.5$ [,+,+[-,+,+][ - , + , + ] (Eq. 3.2.3)
@@ 7.2 Results
In Section 4 we condensed the Roman factorial definition. It is defined as follows:
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 2.2.3)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
Lastly, in Section 6, we constructed these definitions:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\,\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.2: Roman factorial as a rising or falling product
@@
Table Tbl. 6.6.4: Roman factorial as a rising or falling product (generalized)
@@ Acknowledgements
This document was made in  using the online platform Overleaf, compiled in Live 2023. All figures were drawn in Mathematica. ChatGPT 3.5 has assisted in the formatting of this document.
@@ 9.1 Number sets
A set is a mathematical concept representing a collection of distinct items, called elements or members [setwiki] [setfun]. Some sets are so significant in mathematics that they have special names and notations to identify them. These important sets are often denoted using blackboard bold typeface (e.g. $\mathbb{Z}$). Common number sets include the following sets:
@@
$\mathbb{N}\,,\mathbb{Z}^{+}$ Natural numbers 1, 2, 3, 4 $\cdots$
@@
$\mathbb{R}$ Real numbers $\sqrt{2}$, $\pi$, $e$, $\phi$cdots
$\mathbb{I}$ Imaginary numbers $i$, $9.7i$, $-\nicefrac{{i}}{{25}}$cdots
$\mathbb{C}$ Complex numbers $1+i$, $\sqrt{3}-6i$cdots
@@
Note that $\mathbb{C}$ includes all real and imaginary numbers. For instance, the number $2-3i$ is itself a complex number, as well as $5.2i$ and $4$.
@@ 9.2 $\prod$-product
The product operator $\prod$ for a product of a sequence is represented by the capital Greek letter "pi" ($\Pi$), similar to how $\sum$ is used as the summation symbol [product]. For example, the product of the first 6 squares of natural numbers can be expressed as:
@@ 9.3 Falling and rising factorials
The falling factorial (also known as the descending factorial, falling sequential product, or lower factorial [fallrisefact]) is defined as the polynomial
@@
The rising factorial (also known as the Pochhammer function, Pochhammer polynomial, ascending factorial, rising sequential product, or upper factorial [fallrisefact]) is defined as
@@ 9.4 The function $\delta(n)_{f}$
The function $\delta(n)$, defined in Subsection 3.2, is expressed as:
@@
Now, let’s explore the Fourier approximation of the sawtooth function [titchmarsh1986theory]. This approximation, denoted $\{n\}_{f}$ with the subscript $f$ for Fourier, is defined as:
@@
$\lfloor n\rfloor_{f}=n-\{n\}_{f}=n-\frac{1}{2}+\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\;,\enspace\>n\in\mathbb{R}\,.$ (Eq. 9.4.4)
@@
As previously stated in Eq. 3.2.3, $\varepsilon>0$ ensures $\delta(0)=0+\varepsilon>0$, while $\varepsilon<1$ ensures $\delta(-1)=-1+\varepsilon<0$, achieving the desired output pattern [,+,+][-,+,+][ - , + , + ] for all real numbers.
@@
Consequently, $\delta(n)_{f}$ can be used in place of $\delta(n)$ interchangeably. Either definition of Eq. 3.2.3 and Eq. 9.4.6 will suffice, since the F.F. $\theta(n)$ will still have the output pattern $[-1,1,1]$ and $\delta(n)$ does not have any other use.
@@ 9.5 Boolean algebra
Boolean algebra [boolean_algebra] is a branch of algebra named after George Boole, who introduced it in his book The Mathematical Analysis of Logic (1847). It is a formal way of describing logical operations, like logic gates which are used in computing.
@@
A Boolean-valued function [boolean_functions] (sometimes called a predicate or a proposition) is a function whose input is an arbitrary number set and has outputs in a Boolean domain (i.e. a generic two-element set, for example $\{0,1\}$). Those elements are interpreted as logical values, for example, $0=false$ and $1=true$ (a single bit of information).
In Sections 3 and 5, we define a set of Boolean-like functions in order to unify piece-wise definitions of the Roman factorial and its various expansions and generalizations. These functions are termed here foundational functions (F.F.) and are similar to Boolean-valued functions.
F.F. do not entirely belong in Boolean algebra because they have up to 3 different outputs, depending if their input $n$ is greater, less than or equal to 0. Additionally, those outputs are not constrained to be 0 or 1, but also $-1$ and even $n$ in the case of $\Phi(n)$.
However, a few F.F. are considered to be genuine Boolean-valued functions: for instance, $\xi(n)$ is a function that outputs the true/false states of the criterion $n\geq 0$. The function $\Theta(n)$, introduced in Subsection 5.2, can be thought of as an operation between two Boolean-valued functions:
@@ 9.6 The indeterminate form $0^{0}$
The mathematical expression $0^{0}$, known as zero to the power of zero, presents a unique case in mathematics where its definition as either 1 or left undefined depends on the context [0tothe0]. In algebra and combinatorics, it is conventionally defined as $10^{0}=$ 1. However, in mathematical analysis, this expression is often left undefined.
@@
These interpretations converge to $10^{0}=$ 1 in various contexts. We list a few of those contexts here:
  • The binomial theorem $(1+x)^{n}=\sum_{k=0}^{n}\binom{n}{k}x^{k}$ holds true for $x=0$ only if $10^{0}=$ 1.
  • Similarly, in rings of power series, where $x^{0}$ must be defined as 1 for all specializations of $x$, identities like $\frac{1}{1-x}=\sum_{n=0}^{\infty}x^{n}$ and $e^{x}=\sum_{n=0}^{\infty}\frac{x^{n}}{n!}$ hold true for $x=0$ only if $10^{0}=$ 1.
  • In order for the polynomial $x^{0}$ to define a continuous function $f:\mathbb{R}\rightarrow\mathbb{R}$, it is necessary to define $10^{0}=$ 1.
  • In calculus, the power rule $\frac{d}{dx}x^{n}=nx^{n-1}$ is valid for $n=1$ at $x=0$ only if $10^{0}=$ 1.
@@
Generally, $0^{0}$ is defined to be 1 so that many identities and properties stay true for $n=0$. One way of overcoming the anomaly is to use the foundational function $\Psi(n)$ in the base, which is never 0. Thus, an expression like $\Psi(n)^{n}$ is well defined for $n\in\mathbb{R}$. We chose to avoid using $10^{0}=$ 1 in this paper since it wasn’t necessary in any F.F. or generalization.
@@ 9.7 The case of $0!$
The factorial of 00 is defined as $1$, denoted as $0!=1$ [0_factorial]. This definition is justified by several reasons:
@@
[…]
@@
  • This choice aligns with the Gamma function, where $10!=\Gamma(0+1)=$ 1, a necessary condition for the gamma function to be continuous.

### v8

@@ 6.6 $\prod$-product definition step 3: $\theta(n)$
Table Tbl. 6.6.2: Index terms of the products in
Tbl. 6.5.1, highlighting signs and exponents
It is evident that the sign of $k$, as well as the exponent, is represented by the same function. That function has the output pattern [1,1,1][-1,1,1] and a F.F. that matches this requirement is $\theta(n)$:
@@
Table Tbl. 6.6.3: Index terms of the
products in Tbl. 6.5.1, unified
@@
Table Tbl. 6.6.4: Roman factorial as a rising
or falling product (generalized)
This result concludes Section˜6, as well as the first part of this study.
@@
These alternative $\prod$-products would arise if we chose a different approach in Subsections˜6.2 and 6.3, when we chose to multiply the existing products by $n$ instead of altering their upper limits.
@@ 7.1 Summary
  • Section˜2: We analyzed the factorial and presented the Roman factorial.
  • Section˜3: We introduced a set of 5 foundational functions (F.F.).
  • Section˜4: We employed the F.F. shown in the previous section to rewrite the Roman factorial definition concisely, in its original formulation as well as its recursive form.
  • Section˜5: We introduced another set of 5 F.F.
  • Section˜6: We found non-recursive $\prod$-product definitions to describe the values outputted by the Roman factorial. The expressions are either a rising or a falling product, split into two cases each (positive and negative integers). Additionally, we used all F.F. to unify the cases of these definitions, achieving universality.
In Sections˜3 and 5 we defined the following foundational functions:
$\delta(n)$ $=\lfloor n\rfloor+0.5$ $[-,+,+]$ (Eq. 3.2.3)
@@ 7.2 Results
In Section˜4 we condensed the Roman factorial definition. It is defined as follows:
@@
[…]
@@
$\lfloor 0\rceil!=\lfloor-1\rceil!=1\,.$ (Eq. 4.1.1)
@@
$\lfloor 0\rceil!=\lfloor-1\rceil!=1\,.$ (Eq. 4.1.1)
Lastly, in Section˜6, we constructed these definitions:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.2: Roman factorial as a
rising or falling product
@@
Table Tbl. 6.6.4: Roman factorial as a rising or
falling product (generalized)
@@ Acknowledgements
This document was made in  using the online platform Overleaf, compiled in Live 2023. All figures were drawn in Mathematica. ChatGPT 3.5 has assisted in the formatting of this document and not in any research presented here.
@@ 9.1 Number sets
A set is a mathematical concept representing a collection of distinct items, called elements or members [4] [5]. Some sets are so significant in mathematics that they have special names and notations to identify them. These important sets are often denoted using blackboard bold typeface (e.g. $\mathbb{Z}$). Common number sets include the following sets:
@@
$\mathbb{N}\,$$\mathbb{Z}^{+}$ Natural numbers 1, 2, 3, 4 $\cdots$
@@
$\mathbb{R}$ Real numbers $\sqrt{2}$, $\pi$, $e$, $\phi$ $\cdots$
$\mathbb{I}$ Imaginary numbers $i$, $9.7i$, $\nicefrac{{i}}{{25}} \cdots$
$\mathbb{C}$ Complex numbers $1+i$, $\sqrt{3}-6i$ $\cdots$
@@
Note that $\mathbb{C}$ includes all real and imaginary numbers. For instance, the number 23i2-3i is itself a complex number, as well as $5.2i$ and $4$.
@@ 9.2 $\prod$-product
The product operator $\prod$ for a product of a sequence is represented by the capital Greek letter "pi" ($\Pi$), similar to how $\sum$ is used as the summation symbol [6]. For example, the product of the first 6 squares of natural numbers can be expressed as:
@@ 9.3 Falling and rising factorials
The falling factorial (also known as the descending factorial, falling sequential product, or lower factorial [7]) is defined as the polynomial
@@
The rising factorial (also known as the Pochhammer function, Pochhammer polynomial, ascending factorial, rising sequential product, or upper factorial [7]) is defined as
@@ 9.4 The function $\delta(n)_{f}$
The function $\delta(n)$, defined in Subsection˜3.2, is expressed as:
@@
Now, let’s explore the Fourier approximation of the sawtooth function [8]. This approximation, denoted $\{n\}_{f}$ with the subscript $f$ for Fourier, is defined as:
@@
$\lfloor n\rfloor_{f}=n-\{n\}_{f}=n-\frac{1}{2}+\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\>,\enspace\>n\in\mathbb{R}\,.$ (Eq. 9.4.4)
@@
As previously stated in Eq. 3.2.3, $\varepsilon>0$ ensures $\delta(0)=0+\varepsilon>0$, while $\varepsilon<1$ ensures $\delta(-1)=-1+\varepsilon<0$, achieving the desired output pattern [,+,+][-,+,+] for all real numbers.
@@
Consequently, $\delta(n)_{f}$ can be used in place of $\delta(n)$ interchangeably. Either definition of Eq. 3.2.3 and Eq. 9.4.6 will suffice, since the F.F. $\theta(n)$ will still have the output pattern [1,1,1][-1,1,1] and $\delta(n)$ does not have any other use.
@@ 9.5 Boolean algebra
Boolean algebra [9] is a branch of algebra named after George Boole, who introduced it in his book The Mathematical Analysis of Logic (1847). It is a formal way of describing logical operations, like logic gates which are used in computing.
@@
A Boolean-valued function [10] (sometimes called a predicate or a proposition) is a function whose input is an arbitrary number set and has outputs in a Boolean domain (i.e. a generic two-element set, for example $\{0,1\}$). Those elements are interpreted as logical values, for example, $0=false$ and $1=true$ (a single bit of information).
In Sections˜3 and 5, we define a set of Boolean-like functions in order to unify piece-wise definitions of the Roman factorial and its various expansions and generalizations. These functions are termed here foundational functions (F.F.) and are similar to Boolean-valued functions.
F.F. do not entirely belong in Boolean algebra because they have up to 3 different outputs, depending if their input $n$ is greater, less than or equal to 0. Additionally, those outputs are not constrained to be 0 or 1, but also 1-1 and even $n$ in the case of $\Phi(n)$.
However, a few F.F. are considered to be genuine Boolean-valued functions: for instance, $\xi(n)$ is a function that outputs the true/false states of the criterion $n\geq 0$. The function $\Theta(n)$, introduced in Subsection˜5.2, can be thought of as an operation between two Boolean-valued functions:
@@ 9.6 The indeterminate form $0^{0}$
The mathematical expression $0^{0}$, known as zero to the power of zero, presents a unique case in mathematics where its definition as either 1 or left undefined depends on the context [3]. In algebra and combinatorics, it is conventionally defined as $0^{0}=1$. However, in mathematical analysis, this expression is often left undefined.
@@
These interpretations converge to $0^{0}=1$ in various contexts. We list a few of those contexts here:
  • The binomial theorem $(1+x)^{n}=\sum_{k=0}^{n}\binom{n}{k}x^{k}$ holds true for $x=0$ only if $0^{0}=1$.
  • Similarly, in rings of power series, where $x^{0}$ must be defined as 1 for all specializations of $x$, identities like $\frac{1}{1-x}=\sum_{n=0}^{\infty}x^{n}$ and $e^{x}=\sum_{n=0}^{\infty}\frac{x^{n}}{n!}$ hold true for $x=0$ only if $0^{0}=1$.
  • In order for the polynomial $x^{0}$ to define a continuous function $f:\mathbb{R}\rightarrow\mathbb{R}$, it is necessary to define $0^{0}=1$.
  • In calculus, the power rule $\frac{d}{dx}x^{n}=nx^{n-1}$ is valid for $n=1$ at $x=0$ only if $0^{0}=1$.
@@
Generally, $0^{0}$ is defined to be 1 so that many identities and properties stay true for $n=0$. One way of overcoming the anomaly is to use the foundational function $\Psi(n)$ in the base, which is never 0. Thus, an expression like $\Psi(n)^{n}$ is well defined for $n\in\mathbb{R}$. We chose to avoid using $0^{0}=1$ in this paper since it wasn’t necessary in any F.F. or generalization.
@@ 9.7 The case of $0!$
The factorial of 0 is defined as $1$, denoted as $0!=1$ [11]. This definition is justified by several reasons:
@@
  • Defining $0!=1$ ensures that various combinatorial identities hold true for all valid parameter values. For example, the number of ways to select all $n$ elements from a set of $n$ is $\binom{n}{n}=\frac{n!}{n!\,0!}=1$, a binomial coefficient identity valid only if $0!=1$.
@@
  • This choice aligns with the Gamma function, where $0!=\Gamma(0+1)=1$, a necessary condition for the gamma function to be continuous.